### v3

# Kinetically Modified Palatini Inflation Meets ACT Data

Abstract
Abstract: We show that the coexistence of a non-minimal coupling to gravity $f_{\cal R}=1+c_{\cal R}\phi^{n/2}$ with a kinetic mixing of the form $f_{\rm K}=f_{\cal R}^{m}$ – where $n=2$ and $4$ and $0.5\leq m\leq 10$ – reconciles chaotic inflation based on the $\phi^{n}$ potential with the recent ACT results, if we adopt the Palatini formulation of gravity. The attainment of inflation allows for subplanckian inflaton values and energy scales below the cut-off scale of the corresponding effective theory. The model can be also embedded in supergravity by introducing two chiral superfields and a monomial superpotential, linear with respect to the inflaton-accompanying field. Its stabilization is achieved thanks to a compact contribution to the Kähler potential, whose the inflationary part includes an holomorphic logarithmic term and a real one multiplying a shift-symmetric quadratic polynomial term.
PACs numbers: 98.80.Cq, 04.50.Kd, 12.60.Jv, 04.65.+e Published in Phys. Lett. B 868, 139739 (2025)
I Introduction
Working in the context of the metric formulation of gravity – and in the reduced Planck units with $m_{\rm P}=M_{\rm P}/\sqrt{8\pi}=1$ – one can easily verify old ; nmi ; nmiq ; roest that the presence of a non-minimal coupling function
$f_{\cal R}(\phi)=1+c_{\cal R}\phi^{n/2},$ (1)
between the inflaton $\phi$ and the Ricci scalar $\mathcal{R}$, considered in conjunction with a monomial potential of the type
$V_{\rm I}(\phi)=\lambda^{2}\phi^{n}/2^{n/2},$ (2)
provides, at the strong $c_{\cal R}$ limit, an attractor roest towards the (scalar) spectral index
$n_{\rm s}\simeq 1-2/{N_{\star}}=0.965~~\mbox{for}~~{N_{\star}}=55$ (3)
e-foldings with negligible $n_{\rm s}$ running $a_{\rm s}$ and tensor-to-scalar ratio $r\simeq 0.0036$. Focusing on the most well-motivated cases – from Particle Physics point of view –, it would be emphasized that, in the Palatini formulation demir ; attrpal of gravity, the resulting $n_{\rm s}$ in Eq. (3) remains essentially unaltered for $n=4$ whereas it slightly increases for $n=2$ linear ; attrpal . Although perfectly consistent with Bicep2/Keck Array and Planck data plin , the value in Eq. (3) is in tension with the latest Data Release 6 (DR6) from the Atacama Cosmology Telescope (ACT) act ; actin , combined with the cosmic microwave background (CMB measurements by Planck plin and BICEP/Keck (BK) bcp , together with the Dark Energy Spectroscopic Instrument (DESI) Baryon Acoustic Oscillation (BAO) results desi . Indeed, the so-called P-ACT-LB-BK18 data entails actin
$n_{\rm s}=0.974\pm 0.0068,~a_{\rm s}=0.0062\pm 0.0104~~\mbox{and}~~r\leq 0.038,$ (4)
[…]
Following this avenue, we here focus on a variant of non-minimal inflation (nMI) developed nmikr by introducing a suitable non-canonical kinetic mixing $f_{\rm K}(\phi)$ in the inflaton sector. For this reason the term kinetically modified nMI was coined. The key point of our present approach is the adoption of the Palatini rather than the metric formulation of gravity – for a review see Ref. palreview . The consequences of this alteration is twofold with respect to (w.r.t) our findings in Ref. nmikr : The resulting $n_{\rm s}$ increases without an unacceptable elevation of $r$ which can be partially tested in the near future det and the inflationary scale remains below the Ultraviolet (UV) cut-off scale of the effective theory without the introduction of a new parameter ($c_{\rm K}$) – cf. Ref. un1 ; un2 ; riotto ; sibi ; udemir ; un4 ; un3 ; un0 ; un5 . Nonetheless, the kinetic modification is beneficial for Palatini nMI since it assists to obtain subplanckian values for the inflaton field. At last, we present for first time, to our knowledge, a realization of Palatini nMI in the standard Supergravity (SUGRA) following the strategy introduced in Ref. rube ; su11 ; unvr2 .
II Set-up
At the non-SUSY level, our proposal can be formulated in the Jordan frame (JF) where the action of $\phi$ is given by
${\sf S}=\int d^{4}x\sqrt{-\mathfrak{g}}\left(-\frac{f_{\cal R}}{2}\mathcal{R}+\frac{f_{\rm K}}{2}g^{\mu\nu}\partial_{\mu}\phi\partial_{\nu}\phi-V_{\rm I}(\phi)\right).$ (5)
Here $\mathfrak{g}$ is the determinant of the background Friedmann-Robertson-Walker metric, $g^{\mu\nu}$ with signature (+,,,)(+,-,-,-) and we allow for a kinetic mixing through the function $f_{\rm K}(\phi)$ which is conveniently parameterized in terms of $f_{\cal R}$ as follows
$f_{\rm K}(\phi)=f_{\cal R}^{m}~~\mbox{where}~~1/2\leq m\leq 10$ (6)
is an indicative interval for $m$ which assures an ample variation of $f_{\rm K}$. Indeed, any value of $f_{\rm K}$ can be revealed by the parametrization of Eq. (6) selecting conveniently $m=\ln f_{\rm K}/\ln f_{\cal R}$. The considered $m$ margin can be characterized as natural, since the values are of order unity, and allows us to overlaid almost the whole observationally favored region, as we show in Sec. V. Contrary to the models in Ref. martin – with which ours share the same name – we do not violate the quadratic form of the kinetic operator $\dot{}\phi^{2}$ for $\mu=\nu=0$. By performing a conformal transformation nmi , according to which we define the Einstein frame (EF) metric $\widehat{g}_{\mu\nu}=f_{\cal R}\,g_{\mu\nu}$, we can write $\sf S$} in the EF as follows
[…]
where hat is used to denote quantities defined in the EF. Given that $\mathcal{R}$ does not change under the conformal transformation in the Palatini formalism – where the metric and connection are treated as independent variables – the EF canonically normalized field, $\widehat{\phi}$, and potential, $\widehat{V}_{\rm I}$, are defined as follows
$\frac{d\widehat{\phi}}{d\phi}=J=\frac{f_{\rm K}^{1/2}}{f_{\cal R}^{1/2}}=f_{\cal R}^{\frac{m-1}{2}}~~~\mbox{and}~~~\widehat{V}_{\rm I}=\frac{V_{\rm I}}{f_{\cal R}^{2}}=\frac{\lambda^{2}\phi^{n}}{2^{n/2}f_{\cal R}^{2}}.$ (8)
From the expressions in Eq. (8) it is clear that $f_{\rm K}$ through $m$ determines largely the form of $\widehat{\phi}(\phi)$ whereas $f_{\cal R}$ influences exclusively the curvature of $\widehat{V}_{\rm I}$. Therefore, our scheme is expected to be observationally more flexible compared to the traditional (metric or Palatini) nMI where $m=0$.
III Inflationary Dynamics
A period of slow-roll nMI is determined in the EF by the condition
${\footnotesize\sf max}\{\widehat{\epsilon}(\phi),|\widehat{\eta}(\phi)|\}\leq 1,$ (9)
where the slow-roll parameters $\widehat{\epsilon}$ and $\widehat{\eta}$ read
$\widehat{\epsilon}$ $=$ $\left({\widehat{V}_{\rm I,\widehat{\phi}}\over\sqrt{2}\widehat{V}_{\rm I}}\right)^{2}=\frac{n^{2}}{2c_{\cal R}\phi^{2}f_{\cal R}^{1+m}}~~\mbox{and}~~\widehat{\eta}={\widehat{V}_{\rm I,\widehat{\phi}\widehat{\phi}}\over\widehat{V}_{\rm I}}=$ (10)
$=$ $\frac{\widehat{\epsilon}}{2n^{2}}\left(4n(n-1)-n(4+n(m+1))c_{\cal R}\phi^{n/2}\right),$~~
with the symbol $\chi$ as subscript denoting derivation w.r.t the field $\chi$. For the expressions above we employ $J$ in Eq. (8), without expressing explicitly $\widehat{V}_{\rm I}$ in terms of $\widehat{\phi}$. For some $m$ values, actually, this is not doable due to the complicate form of the $\widehat{\phi}(\phi)$ function which is the following
$\widehat{\phi}$ $=$ $\phi f_{\cal R}^{\frac{m+1}{2}}\;{}_{2}F_{1}\left(1,\frac{m}{2}+\frac{2}{n}+\frac{1}{2};1+\frac{2}{n};-c_{\cal R}\phi_{\star}^{n/2}\right)$ (11)
$\simeq$ $\frac{4c_{\cal R}^{(m-1)/2}\phi^{1+n(m-1)/4}}{4+n(m-1)}~~\mbox{for}~~c_{\cal R}\gg 1.$
Here $_{2}F_{1}$ is the Gauss hypergeometric function wolfram . Note that our expression reduces to that given in Ref. attrpal for $m=0$ whereas for $m=1$ we obtain a case with $\widehat{\phi}=\phi$, recently mentioned for $n=4$ in Ref. un5 . Eq. (9) is saturated at the maximal $\phi$ value, $\phi_{\rm f}$, from the following two values
$\left(\phi_{1\rm f},\phi_{2\rm f}\right)\simeq\Bigg{(}\left(\frac{n^{2}}{2c_{\cal R}^{m+1}}\right)^{\frac{2}{4+n(m+1)}},\left(\frac{n(n+1)}{c_{\cal R}^{1+m}}\right)^{\frac{2}{4+n(m+1)}}\Bigg{)}$ (12a)
which satisfy for $\phi\ll 1$ the equations
$\widehat{\epsilon}\left(\phi_{1\rm f}\right)=|\widehat{\eta}\left(\phi_{2\rm f}\right)|\simeq 1.$ (12b)
Table 1: Inflationary predictions for $n=2$ and $4$.
$n=2$ $n=4$
$n_{\rm s}$ $1-(m+3)\big{(}1+1/(2(m+2)c_{\cal R}^{2}{N_{\star}})^{1/(m+2)}\big{)}/(m+2){N_{\star}}$ $1-(m+2)\big{(}1+1/(8(m+1)c_{\cal R}{N_{\star}})^{1/(m+1)}\big{)}/(m+1){N_{\star}}$
$r$ $16/(m+2){N_{\star}}\big{(}2c_{\cal R}^{2}(m+2){N_{\star}})\big{)}^{1/(m+2)}$ $16/(m+1){N_{\star}}\big{(}8c_{\cal R}(m+1){N_{\star}})\big{)}^{1/(m+1)}$
$a_{\rm s}$ $(m+3)/(m+2)N^{2}_{\star}-{\cal O}\big{(}N^{-2}_{\star}(c_{\cal R}^{2}{N_{\star}})^{-1/(m+2)}\big{)}$ $(m+2)/(m+1)N^{2}_{\star}-{\cal O}\big{(}N^{-2}_{\star}(c_{\cal R}{N_{\star}})^{-1/(m+1)}\big{)}$
The number of e-foldings $N_{\star}$} that the scale $k_{\star}=0.05/{\rm Mpc}$ experiences during this nMI and the amplitude $A_{\rm s}$ of the power spectrum of the curvature perturbations generated by $\phi$ can be computed using the standard formulae
${N_{\star}}=\int_{\widehat{\phi}_{\rm f}}^{\widehat{\phi}_{\star}}d\widehat{\phi}\frac{\widehat{V}_{\rm I}}{\widehat{V}_{\rm I,\widehat{\phi}}}~~\mbox{and}~~A_{\rm s}^{1/2}=\frac{1}{2\sqrt{3}\,\pi}\;\frac{\widehat{V}_{\rm I}^{3/2}(\widehat{\phi}_{\star})}{|\widehat{V}_{\rm I,\widehat{\phi}}(\widehat{\phi}_{\star})|}$, (13)
where $\phi_{\star}~[\widehat{\phi}_{\star}$] is the value of $\phi~[\widehat{\phi}$] when $k_{\star}$ crosses the inflationary horizon. Since $\phi_{\star}\gg\phi_{\rm f}$, from Eq. (13) we find
${N_{\star}}$ $=$ $\frac{\phi_{\star}^{2}}{2n}\;{}_{2}F_{1}\left(-m,4/n;1+4/n;-c_{\cal R}\phi_{\star}^{n/2}\right)$ (14)
$\simeq$ $\frac{2c_{\cal R}^{m}\phi_{\star}^{(nm+4)/2}}{n(nm+4)}~~\mbox{for}~~c_{\cal R}\gg 1.$
For $c_{\cal R}\gg 1$, we are able to solve Eq. (14) w.r.t $\phi_{\star}$ and pursue our analytical approach. Indeed, we obtain
$\phi_{\star}\simeq\left(n\left(nm+4\right){N_{\star}}/2c_{\cal R}^{m}\right)^{2/(4+nm)}.$ (15)
It is clear that there is a lower bound on $c_{\cal R}$, above which $\phi_{\star}\leq 1$ and so, our proposal can be stabilized against corrections from quantum gravity. I.e.,
$\phi_{\star}\leq 1~~\Rightarrow~~c_{\cal R}\geq\left(n\left(nm+4\right){N_{\star}}/2\right)^{1/m}.$ (16)
Note that this bound can be imposed only for $m\neq 0$, as advocated in Sec. I, and not within the pure Palatini nMI. Moreover, the minimal $c_{\cal R}$ value decreases as $m$ increases and so problems with the pertubative unitarity – see below – become less acute.
From the rightmost relation in Eq. (13) we can also derive a constraint between $\lambda$ and $c_{\cal R}$ solving w.r.t $\lambda$. I.e.
$\lambda\simeq 2^{\frac{n+4}{4}}(3A_{\rm s})^{\frac{1}{2}}\pi nc_{\cal R}^{\frac{nm+2}{nm+4}}(2/n(nm+4){N_{\star}})^{\frac{4+n(m+1)}{2(nm+4)}}.$ (17)
For $m=0$ we reproduce the $\lambda-c_{\cal R}$ relation obtained in the pure Palatini nMI which differs slightly from that within the metric nMI – cf. Ref. palreview .
The inflationary observables are found from the relations
$n_{\rm s}=\>1-6\widehat{\epsilon}_{\star}\ +\ 2\widehat{\eta}_{\star},~~r=16\widehat{\epsilon}_{\star}$, (18a)
$a_{\rm s}=\>2\left(4\widehat{\eta}_{\star}^{2}-(n_{\rm s}-1)^{2}\right)/3-2\widehat{\xi}_{\star}$, (18b)
where the variables with subscript $\star$ are evaluated at $\phi=\phi_{\star}$ and $\widehat{\xi}={\widehat{V}_{\rm I,\widehat{\phi}}\widehat{V}_{\rm I,\widehat{\phi}\widehat{\phi}\widehat{\phi}}/\widehat{V}^{2}_{\rm I}}$. For $c_{\cal R}\gg 1$ we find the following general – i.e., for any $n$ and $m$ – expressions
$n_{\rm s}=1-\frac{1}{{N_{\star}}}-\frac{n}{(nm+4){N_{\star}}}$
$-\left(\frac{2^{n}}{n^{n}c_{\cal R}^{4}}\right)^{\frac{1}{nm+4}}\frac{n(m+1)+4}{((nm+4){N_{\star}})^{1+\frac{n}{nm+4}}};$ (19a)
$r=16\left(\frac{2^{n}n^{4+n(m-1)}}{c_{\cal R}^{4}((nm+4){N_{\star}})^{4+n(m+1)}}\right)^{\frac{1}{nm+4}}.$ (19b)
As regards $a_{\rm s}$, specific expressions for $n=2$ and $4$ are given in Table 1 together with the corresponding outputs of Eqs. (19a) and (19b). Taking the limit $c_{\cal R}\rightarrow\infty$ for $m=0$, Eq. (19a) is in agreement with Ref. linear ; attrpal
$n_{\rm s}=1-1/{N_{\star}}-n/4{N_{\star}},$ (20)
if we take into account the appropriate redefinition of $n$. Therefore, for $n=4$ we arrive again at Eq. (3) whereas for $n=2$ we obtain $n_{\rm s}=1-3/2{N_{\star}}$ which reaches Eq. (4) for ${N_{\star}}\simeq 60$. The same is true for $n=4$ and $m=1$ – cf. Ref. un5 . From the results of Table 1 we can clearly infer that increasing $m$, $n_{\rm s}$ increases whereas $r$ remains moderately suppressed due to the factor $c_{\cal R}^{-4/(nm+4)}$.
IV Effective Cut-Off Scale
It is well known udemir that the absence of an extra term in the leftmost expression in Eq. (8) – cf. Ref. nmikr – deliberates the Palatini nMI for $n=4$ from the problem with the perturbative unitarity which puzzles the metric nMI un1 ; un2 ; riotto – for another attitude on this issue see Ref. sibi . We show here that this nice feature of Palatini nMI insists even in our cases.
To this end we follow the most restrictive approach analyzing the small-field behavior of our models in the EF – cf. Ref. riotto ; un1 ; un2 ; nmikr . We focus on the second term in the right-hand side of Eq. (7) – or Eq. (28a), see Sec. VI below – for $\mu=\nu=0$ and we expand it about $\langle{\phi}\rangle=0$ in terms of $\widehat{\phi}$ – see Eq. (8) – which coincides with $\phi$ since $\langle{J}\rangle=1$. Our result for some $m$ values can be written as
$J^{2}\dot{\phi}^{2}=\dot{}\phi^{2}\begin{cases}1&\mbox{for}~~m=1,\\ 1+c_{\cal R}\phi^{n/2}&\mbox{for}~~m=2,\\ \left(1+2c_{\cal R}\phi^{n/2}+c_{\cal R}^{2}\phi^{n}+\cdots\right)&\mbox{for}~~m=3.\end{cases}$
Similar expressions can be obtained for the other $m$’s too. Expanding similarly $\widehat{V}_{\rm I}$, see Eq. (8), in terms of $\phi$ we have
$\widehat{V}_{\rm I}=\frac{\lambda^{2}\phi^{n}}{2^{n/2}}\left(1-2c_{\cal R}\phi^{\frac{n}{2}}+3c_{\cal R}^{2}\phi^{n}-4c_{\cal R}^{3}\phi^{\frac{3n}{2}}+\cdots\right)$
independently of $m$. If we reinstall $m_{\rm P}$ in the expressions above, so that we obtain dimensionless quantities in the bracket, and introduce $c_{\cal R}$ in the denominators, we can conclude that the UV cut-off scale is
$\Lambda_{\rm UV}=m_{\rm P}/c_{\cal R}^{2/n}~~\mbox{for any $m$.}$ (21)
Our result is in accordance with Ref. udemir ; un0 ; un3 ; un4 for the $n=4$ and $m=0$ case. Consistency with the effective theory entails
$\widehat{V}_{\rm I0}^{1/4}\leq\Lambda_{\rm UV}~~\mbox{with}~~\widehat{V}_{\rm I0}\simeq\lambda^{2}m_{\rm P}^{4}/2^{n/2}c_{\cal R}^{2}.$ (22)
If we take into account Eq. (17), the condition above yields an upper bound on $c_{\cal R}$ for every selected $n$ and $m$. As we show in Sec. V, there is sizable parameter space consistent with this requirement – even for the most restrictive case with $n=2$. The validity of Eq. (22) is of crucial importance for the viability of the effective theory riotto , independently from the hierarchy between $\phi_{\star}$ and $\Lambda_{\rm UV}$, since it is widely believed that dangerous loop-corrections depend on the energy scale and not on the field values. Therefore, the fact that in our models we obtain $\phi_{\star}\gg\Lambda_{\rm UV}$ – which causes, in principle, concerns regarding corrections from non-renormalizable terms associated with $\Lambda_{\rm UV}$ – does not invalidate our proposal. In addition, a more elaborated derivation of $\Lambda_{\rm UV}$ un0 for $n=4$ and $m=0$ reveals the absence of any relevant problem.
Fig. 1: Allowed curves in the $n_{\rm s}-r$ plane for $n=2$ (left panel) and $4$ (right panel) and various $m$’s indicated on the curves. We take ${N_{\star}}=50$ (dashed lines) or ${N_{\star}}=60$ (solid lines) and minimal (thick lines) or maximal (thin lines) $c_{\cal R}$ values. The marginalized joint $68\%$ $95\%$] c.l. regions from P-ACT-LB-BK18 data are depicted by the dark [light] shaded contours.
V Numerical Results
The approximate analytical estimates in Sec. III – which are accurate enough for $c_{\cal R}\gg 1$ – can be verified and extended to all possible $c_{\cal R}$ values numerically. Namely, confronting the quantities in Eq. (13) with the observational requirements act
$50\lesssim{N_{\star}}\lesssim 60~~\mbox{and}~~A_{\rm s}^{1/2}\simeq 4.618\cdot 10^{-5},$ (23)
we can restrict $\lambda$ and $\phi_{\star}$ and compute the model predictions via Eqs. (18a) and (18b), for any selected $c_{\cal R}$, $m$ and $n$. Taking the $N_{\star}$} range above we assumed that nMI is followed in turn by a oscillatory phase with mean equation-of-state parameter $w_{\rm rh}$, radiation and matter domination, with variation plin ; rehpal ; rehpal1 ; reh1 ; reh2 of $w_{\rm rh}$ and the reheating temperature $T_{\rm rh}$ in the following ranges respectively
$0\lesssim w_{\rm rh}\lesssim 1/3~~\mbox{and}~~1~{\rm TeV}\lesssim T_{\rm rh}\lesssim 1~{\rm EeV},$ (24)
with units reinstalled. The $c_{\cal R}$ values can be further constrained by the saturation of Eqs. (16) and (22), from which the first one gives the lowest $c_{\cal R}$ value (for $\phi_{\star}=1$) whereas the second yields the largest one.
The outputs of our numerical approach, encoded as lines in the $n_{\rm s}-r$ plane, are compared against the P-ACT-LB-BK18 data actin in Fig. 1 for $n=2$ (left panel) and $n=4$ (right panel). The variation of $m$ is shown along each line. In each plot we depict solid lines for ${N_{\star}}=60$ and dashed lines for ${N_{\star}}=50$. We also display two thick lines corresponding to the minimal possible $c_{\cal R}$ and two thin ones originating from the maximal $c_{\cal R}$. E.g., for $m=1$ and ${N_{\star}}=50$ or $60$ we obtain correspondingly
$302\lesssim c_{\cal R}\lesssim 625$ or $362\lesssim c_{\cal R}\lesssim 670~~\mbox{for}~~n=2~~\mbox{and}$
$8\lesssim\frac{c_{\cal R}}{100}\lesssim 3.5\cdot 10^{3}$ or $9.7\lesssim\frac{c_{\cal R}}{100}\lesssim 4.2\cdot 10^{3}~~\mbox{for}~~n=4.$
In the domain of large $c_{\cal R}$ we assure that the perturbative bound $\lambda=3.5$ is not violated. Decreasing $m$, the needed $c_{\cal R}$ values increase and so Eq. (22) becomes more relevant. For $n=2$ and $m=0.88$ and $0.87$ the lower and upper bounds on $c_{\cal R}$ from Eqs. (16) and (22) respectively coincide and so the the two (thick and thin) dashed and solid lines intersect. On the other hand, the graph for $n=4$ is extended to lower $m$ values and such an intersection does not occur for the range of $m$ in Eq. (6). E.g., along the dashed thick lines, with decreasing $m$, we obtain
$0.005\lesssim\widehat{V}_{\rm I0}^{1/4}/\Lambda_{\rm UV}\lesssim 1~[0.07]~~\mbox{for $n=2$ [$n=4$]}.$ (25)
On the other hand, as $m$ increases $r$ increases slowly and slowly due to decrease of $c_{\cal R}$ – see Eq. (19b) – and so the upper bound in Eq. (4) is not saturated. Therefore, no realistic upper bound on $m$ can be found.
Comparing the resulting corridor between the two thick (or thin) lines in the two panels of Fig. 1 we remark that the allowed region for $n=4$ overlaps better with 1-$\sigma$ range of Eq. (4). Indeed, as $n$ increases above $2$ the resulting corridor moves to the left. E.g., fixing $m=\phi_{\star}=1$ we obtain
$0.969\lesssim n_{\rm s}\lesssim 0.974$ for $n=6,$ (26a)
$0.967\lesssim n_{\rm s}\lesssim 0.973$ for $n=8,$ (26b)
$0.966\lesssim n_{\rm s}\lesssim 0.971$ for $n=10,$ (26c)
with $r$ decreasing as $n$ increases. Therefore, the considered here $n$ values offer the best prospects for detecting primordial gravitational waves in future experiments det .
Note, finally, that along the curves depicted in Fig. 1 the obtained $a_{\rm s}$ remains negligibly small and therefore it lies within the margin in Eq. (4). E.g., along the dashed [solid] thick lines, it is $a_{\rm s}\simeq 5.3\cdot 10^{-4}~[a_{\rm s}\simeq 3.6\cdot 10^{-4}]$ for $n=2$ whereas for $n=4$, we obtain
$5.2\lesssim-a_{\rm s}/10^{-4}\lesssim 6.3~[3.7\lesssim-a_{\rm s}/10^{-4}\lesssim 4.3]$ (27)
with decreasing $m$. These results are of the same order of magnitude with those obtained within conventional (metric or Palatini) nMI.
VI Supergravity Embedding
Starting from the EF action in SUGRA we attempt to reproduce the ingredients of our model in Eq. (8) without caring if the complete JF SUGRA action is consistent with Eq. (5). This task is not realized up to now, to our knowledge, for the Palatini formalism.
Adopting the strategy of Ref. rube ; su11 ; unvr2 we consider two gauge-singlet chiral superfields, i.e., $z^{\alpha}=\Phi,S$, with $\Phi$ (${\alpha}=1$) and $S$ (${\alpha}=2)$ being the inflaton and a “stabilizer” field respectively. The EF action for $z^{\alpha}$’s within SUGRA can be written as
$\hskip-5.69054pt{\sf S}=\int d^{4}x\sqrt{-\widehat{\mathfrak{g}}}\left(-\frac{1}{2}\widehat{\mathcal{R}}+K_{{\alpha}{\bar{\beta}}}\widehat{g}^{\mu\nu}\partial_{\mu}z^{\alpha}\partial_{\nu}z^{*{\bar{\beta}}}-\widehat{V}_{\rm F}\right)$ (28a)
where summation is taken over the scalar fields $z^{\alpha}$, star () denotes complex conjugation, $K$ is the Kähler potential with $K_{{\alpha}{\bar{\beta}}}=K_{,z^{\alpha}z^{*{\bar{\beta}}}}$ and $K^{{\alpha}{\bar{\beta}}}K_{{\bar{\beta}}\gamma}=\delta^{\alpha}_{\gamma}$. Also $\widehat{V}_{\rm F}$ is the EF F–term SUGRA potential given by
$\widehat{V}_{\rm F}=e^{K}\left(K^{{\alpha}{\bar{\beta}}}(D_{\alpha}W)(D^{*}_{\bar{\beta}}W^{*})-3{|W|^{2}}\right),$ (28b)
where $D_{\alpha}W=W_{,z^{\alpha}}+K_{,z^{\alpha}}W$ with $W$ being the superpotential. Therefore, the desired SUGRA embedding of our models requires the determination of $K$ and $W$.
Our task can be facilitated, if we express $\Phi$ and $S$ according to the parametrization
$\Phi=\>{\phi\,e^{i{\theta}}}/{\sqrt{2}}\>\>\>\mbox{and}\>\>\>S=\>(s+i\bar{s})/\sqrt{2}\,$, (29)
and determine the inflationary track by the constraints
$\langle{S}\rangle_{\rm I}=\langle{\Phi-\Phi^{*}}\rangle_{\rm I}=0,~\mbox{or}~~\langle{s}\rangle_{\rm I}=\langle{\bar{s}}\rangle_{\rm I}=\langle{{\theta}}\rangle_{\rm I}=0,$ (30)
where the symbol $\langle{Q}\rangle_{\rm I}$ stands for the value of a quantity $Q$ during nMI. $V_{\rm I}$ in Eq. (2) can be produced, in the flat limit, by
$W=\lambda S\Phi^{n/2}~~\mbox{since}~~\langle{|W_{,S}|^{2}}\rangle_{\rm I}=V_{\rm I}.$ (31)
The form of $W$ can be uniquely determined if we impose an $R$ symmetry, under which $S$ and $\Phi$ have charges $1$ and 0, and a global $U(1)$ symmetry with assigned charges to $S$ and $\Phi$ 1-1 and $2/n$. The latter is violated, though, in the proposed $K$ which is judiciously chosen so that $J$ and $\widehat{V}_{\rm I}$ in Eq. (8) are reproduced. More specifically, $\widehat{V}_{\rm I}$ has to be derived from the only surviving term in Eq. (28b) along the track in Eq. (30), which is
$\langle{\widehat{V}_{\rm F}}\rangle_{\rm I}=\langle{e^{K}K^{SS^{*}}\,|W_{,S}|^{2}}\rangle_{\rm I}\,.$ (32)
The proposed $K$ includes two contributions without mixing between $\Phi$ and $S$, i.e.,
$K=K_{1}+K_{2},$ (33)
from which $K_{2}$ successfully stabilizes $S$ along the path in Eq. (30) without invoking higher order terms – cf. Ref. nmikr . We adopt the form su11
$K_{2}=N_{2}\ln\left(1+{|S|^{2}/N_{2}}\right)~~\mbox{with}~~0<N_{2}<6$ (34)
which parameterizes su11 the compact manifold $SU(2)/U(1)$ with curvature $2/N_{2}$. On the other hand, $K_{1}$ depends on $\Phi$ and has to incorporate $f_{\cal R}$ and $f_{\rm K}$ in Eqs. (1) and (6). To this end we introduce the functions
$F_{\cal R}(\Phi)=1+2^{\frac{n}{4}}c_{\cal R}\Phi^{\frac{n}{2}}~~\mbox{and}~~F_{\rm sh}=-\frac{1}{2}(\Phi-\Phi^{*})^{2},$ (35)
from which $F_{\cal R}$ is holomorphic, reducing to $f_{\cal R}$ along the path in Eq. (30), whereas $F_{\rm sh}$ is real, assisting us to incorporate the non-canonical kinetic mixing in Eq. (8) – cf. Ref. nmikr . In particular,
$\langle{F_{\cal R}}\rangle_{\rm I}=f_{\cal R},~\langle{F_{\rm sh}}\rangle_{\rm I}=0~~\mbox{with}~~F_{\rm sh,\Phi\Phi^{*}}=1.$ (36)
Therefore, $F_{\cal R}$ can be used to derive the denominator of $\widehat{V}_{\rm I}$ in Eq. (8) whereas the shift-symmetric $F_{\rm sh}$, multiplied by a real function of $\Phi$, lets intact $\langle{\widehat{V}_{\rm F}}\rangle_{\rm I}$ but it may contribute to the normalization of $\phi$. The relevant functions $f_{\rm K}$ and $J$ can be found in terms of $F_{\cal R}$ during nMI as follows
$f_{\rm K}=\langle{(F_{\cal R}+F_{\cal R}^{*})/2}\rangle_{\rm I}^{m}~~\mbox{and}~~J^{2}=\langle{(F_{\cal R}+F_{\cal R}^{*})/2}\rangle_{\rm I}^{m-1}.$ (37)
Employing the ingredients above, a simple, suitable form of $K_{1}$ in Eq. (33) is
$K_{1}=-\ln F_{\cal R}-\ln F_{\cal R}^{*}+\left(\frac{F_{\cal R}+F_{\cal R}^{*}}{2}\right)^{m-1}F_{\rm sh}.$ (38)
It is worth noticing that contrary to the embedding of metric nMI in SUGRA – cf. Ref. nmikr –, $F_{\cal R}$ and $F_{\cal R}^{*}$ enter two different terms in the right-hand side of Eq. (38) and so they give zero contribution into $K_{\Phi\Phi^{*}}$. Note that for $m=1$, $F_{\rm sh}$ and $F_{\cal R}$ are totally decoupled. If, in addition, we take $n=4$ then $K$ is totaly quadratic. Therefore, this is the most well-motivated case from the point of view of SUGRA. Mild (of order $0.1$) variation of the prefactors of the logarithms is expected to have some impact on the observational results – cf. Ref. nmiq ; jhep – at the cost of some tuning, though, that we avoid in our current investigation.
The appropriateness of $K$ in Eqs. (33), (34) and (38) can be verified if we insert it into Eq. (32) and prove
$\langle{K_{\Phi\Phi^{*}}}\rangle_{\rm I}=J^{2}~~\mbox{and}~~\langle{\widehat{V}_{\rm F}}\rangle_{\rm I}=\widehat{V}_{\rm I},$ (39)
where $J$ and $\widehat{V}_{\rm I}$ are given in Eq. (8). Indeed, the first of the expressions above is derived if we take into account that
$\langle{K_{\Phi\Phi^{*}}}\rangle_{\rm I}=\langle{K_{1\Phi\Phi^{*}}}\rangle_{\rm I}=J^{2}\langle{F_{\rm sh,\Phi\Phi^{*}}}\rangle_{\rm I}=J^{2}$ (40a)
– see Eq. (36) – whereas the second one, if we consider Eq. (31) and notice that
$\langle{e^{K}}\rangle_{\rm I}=f_{\cal R}^{-2}~~\mbox{and}~~\langle{K_{SS^{*}}}\rangle_{\rm I}=\langle{K^{SS^{*}}}\rangle_{\rm I}=1.$ (40b)
Consequently, both expressions in Eq. (8) can be recovered from the proposed $W$ and $K$ in Eqs. (31) and (33) via Eq. (39).
Table 2: Mass spectrum along the path in Eq. (30).
Fields Eingestates Masses Squared
Scalars
11 real $\widehat{\theta}$ $(6+3n^{2}c_{\cal R}/2\phi^{2-n/2}f_{\cal R}^{m+1})\widehat{H}_{\rm I}^{2}$
22 real $\widehat{s},\widehat{\bar{s}}$ $(6/N_{2}+3n^{2}/2\phi^{2}f_{\cal R}^{m+1})\widehat{H}_{\rm I}^{2}$
Spinors
22 Weyl $({\widehat{\psi}_{S}\pm\widehat{\psi}_{\Phi})/\sqrt{2}}$~~ $3n^{2}\widehat{H}_{\rm I}^{2}/2\phi^{2}f_{\cal R}^{m+1}$
To consolidate the SUGRA embedding of our models we further verify that the configuration in Eq. (30) is stable w.r.t the excitations of the non-inflaton fields which are canonically normalized fields by the relations
$\widehat{{\theta}}=J{\theta}\phi~~\mbox{and}~~(\widehat{s},\widehat{\bar{s}})={(s,\bar{s})}.$ (41)
In particular, we find the expressions of the masses squared $\widehat{m}^{2}_{\chi^{\alpha}}$ (with $\chi^{\alpha}={\theta}$ and $s$) arranged in Table 2. These expressions assist us to appreciate the role of $N_{2}$ with $0<N_{2}<6$ in retaining positive $\widehat{m}^{2}_{s}$ – in practise we use $N_{2}=1$. Also we confirm that $\widehat{m}^{2}_{\chi^{\alpha}}\gg\widehat{H}_{\rm I}^{2}=\widehat{V}_{\rm I0}/3$ for $\phi_{\rm f}\leq\phi\leq\phi_{\star}$. In Table 2 we display the masses $\widehat{m}^{2}_{\psi^{\pm}}$ of the corresponding fermions too. We define $\widehat{\psi}_{S}=\sqrt{K_{SS^{*}}}\psi_{S}$ and $\widehat{\psi}_{\Phi}=\sqrt{K_{\Phi\Phi^{*}}}\psi_{\Phi}$ where $\psi_{\Phi}$ and $\psi_{S}$ are the Weyl spinors associated with $S$ and $\Phi$ respectively.
Inserting the derived mass spectrum in the well-known Coleman-Weinberg formula, we can find the one-loop radiative corrections, $\Delta\widehat{V}_{\rm I}$ to $\widehat{V}_{\rm I}$ – cf. Ref. nmiq ; jhep . It can be verified that our results are immune from $\Delta\widehat{V}_{\rm I}$, provided that the renormalization group mass scale $\Lambda_{\rm CW}$, is determined by requiring $\Delta\widehat{V}_{\rm I}(\phi_{\star})=0$ or $\Delta\widehat{V}_{\rm I}(\phi_{\rm f})=0$. E.g., imposing the first of these conditions we find $\Lambda_{\rm CW}\simeq(3.2-41)\cdot 10^{-6}$ for $n=2$ and $\Lambda_{\rm CW}\simeq(7.75-39)\cdot 10^{-6}$ for $n=4$ along the thick dashed lines of Fig. 1. Under these circumstances, our results in the SUGRA set-up can be reproduced by using exclusively the ingredients of Eq. (8) as in the non-SUSY set-up.
VII Conclusions
Prompted by the P-ACT-LB-BK18 data actin which is just marginally compatible with the conventional (metric or Palatini) nMI nmi ; roest ; attrpal , we proposed a variant within Palatini gravity which assures comfortable consistency with the current observations.
Namely, working along the lines of Ref. nmikr , we considered a non-canonical kinetic mixing, Eq. (6), – involving the exponent $m$ – in the inflaton sector, apart from the non-minimal coupling to gravity, Eq. (1), which is associated with the potential in Eq. (2). Confining $m$ to the range (0.87[0.5]10)(0.87~[0.5]-10) for $n=2$ [$n=4$] we achieved observational predictions which offer a nice covering of the present data on $n_{\rm s}$ – its compatibility especially with the $n=4$ case for low $m$ values is really impressive, as shown in Fig. 1. Part of the resulting $r$ values there could be accessible in the near future. Our solutions can be attained with subplanckian values of the inflaton, requiring large $c_{\cal R}$’s, and without causing serious problems with the perturbative unitarity since the unitarity-violation scale remains well above the inflationary one. It is gratifying, in addition, that a sizable fraction of the allowed parameter space of our models (with $c_{\cal R}\gg 1$) can be studied analytically.
Our setting can be elegantly implemented in SUGRA too, employing the super- and Kähler potentials given in Eqs. (31) and (33) with individual contributions in Eqs. (34) and (38). The non-minimal coupling of the inflaton to gravity is generated by an holomorphic logarithmic contribution into $K$ whereas its kinetic mixing is expressed by a real function multiplying a shift-symmetric quadratic polynomial term. Our construction can be extended for a gauge non-singlet Higgs superfield as done in Ref. nmhkr ; jhep for the metric formulation.
One possible shortcoming of our proposal is that it employs one extra parameter ($m$) compared to the original model of nMI reducing, thereby, its predictability. As another consequence, however, the enhancement of $n_{\rm s}$ for $n=2$ with $m>0$ may be more drastic than that presented in Ref. linear ; actlinde . Moreover, within our scheme, the reheating phase is not constrained as, e.g., in Ref. rha ; rhc ; act5 ; r2drees ; rhb and corrections to the inflationary potential from other sectors of the theory are not required as, e.g., in Ref. act6 ; oxf ; warm .
Acknowledgments
I would like to thank S. Ketov for interesting discussions.